27
Apr
2007
16:31

This Friday (after midnight) the L2TP machine is being reloaded. This should not result in any problems. If on Saturday there is no connectivity on one of your lines please reboot the appropriate router.

On Sunday, one of the 2 core routers is being replaced. This should have no impact as the network is designed to operate on just one of them should the other fail. If your connection is lost for some reason during this operation *do not* reboot your router.

If you have any problems please call our support number +44 20 7517 4902

Posted by admin

23
Apr
2007
00:05

There has been a spate of short outages for ADSL customers this evening. An earlier DDoS attack had caused a core router to revert to running an old version of software, which was unstable under certain specific conditions which unfortunately occurred this evening. It has now been updated and service should now be stable. If any lines remain out of service, power-cycling the ADSL router should restore them.

Apologies for the inconvenience.

Posted by admin

04
Apr
2007
15:21

We have just experienced another DDoS attack, which once again resulted in loss of service for most customers. Filtering has been employed and service has returned to normal. Apologies for the inconvenience

Posted by admin

Current Status

    GoodNo current issues

Status Archive